For its three investment centers, Gerrard Company accumulates the following data:

I

II

III

Sales $2,051,000 $4,078,000 $4,035,000
Controllable margin 1,256,000 2,018,250 4,012,800
Average operating assets 5,024,000 8,073,000 12,160,000

Compute the return on investment (ROI) for each center.

I

II

III

The return on investment

%

%

%
Answer #1

ROI = Controllable margin/Average operating assets

.

I = 1,256,000/5,024,000 = 25%

II = 2,018,250/8,073,000 = 25%

III = 4,012,800/12,160,000 = 33%